OpenStudy (anonymous):

Please help with exponential and log functions ! I'm new here... 17. Solve in 2 + in x = 5. Round to the nearest thousandth, if necessary.

Parth (parthkohli):

You mean ln(2) + ln(x)?

OpenStudy (anonymous):

i suppose, i just wrote it how it was on the lesson.

Parth (parthkohli):

Well, that is a lowercase L. Anyway. Do you know what ln(a) + ln(b) is?

OpenStudy (anonymous):

no :( i'm really bad at math

Parth (parthkohli):

It is an identity that ln(a) + ln(b) = ln(ab).

OpenStudy (anonymous):

ohhh it's an l lol i feel stupid

OpenStudy (anonymous):

and okay i see

Parth (parthkohli):

So ln(2) + ln(x) can be written as ln(2x).

OpenStudy (anonymous):

i understand so far

Parth (parthkohli):

Can you write 5 in terms of ln?

OpenStudy (anonymous):

umm.. would it be ln(5) ?

Parth (parthkohli):

Nope. Do you know what ln is?

OpenStudy (anonymous):

natural log?

Parth (parthkohli):

Yeah. For example, what is ln(e^2)?

OpenStudy (anonymous):

2?

Parth (parthkohli):

That is correct! So can you get 5 in terms of something's ln?

OpenStudy (anonymous):

ln(e^5) maybe?

Parth (parthkohli):

Yay!

Parth (parthkohli):

ln(2x) = ln(e^5)

Parth (parthkohli):

You should be able to do it now

OpenStudy (anonymous):

i understand now! yay lol thanks!

Parth (parthkohli):

Yeah, no problem!